OpenStudy (anonymous):

Your medical bill is $2345. Your health insurance covers 70% after a $120 deductible. What amount of the bill will you pay?

OpenStudy (anonymous):

So this time it says the insurance will cover 70% AFTER the deductible, so 2345-120=2225. Then it says how much will you be paying so 100-70=30. So again do proportions, cross-multiply and solve for x.\[\frac{ 30 }{ 100 }=\frac{ x }{ 2225 }\]If you need help with a another problem, just put it here.

OpenStudy (anonymous):

Your annual health insurance premium is $3000. Your business pays 65%. If you are paid weekly, what is the paycheck deduction?

OpenStudy (anonymous):

Ok same thing, first, you'll be paying 35% since your business pays 65%. So do proportions cross-multiply and find x. \[\frac{ percentage }{ total percentage }=\frac{ amount }{ totalamount }\]So \[\frac{ 35 }{ 100 }=\frac{ x }{ 3000 }\]Once you find x this time its asking how much will you be paying weekly. Since the premium was annually (yearly), divide x by 52. 52 weeks=1 year.